A taxi left Los Angeles at 7:50 am travelling 50 kph. At 9 am, an express train travelling 110 kph left the same spot travelled by the taxi. At what time did the train overtake the taxi
Found 2 solutions by palanisamy, Alan3354:
Answer by palanisamy(496)

A taxi left Los Angeles at 7:50 am travelling 50 kph. At 9 am, an express train travelling 110 kph left the same spot travelled by the taxi.
Let the taxi and the train meet x hours after 9AM.
Distance travelled by the taxi = (x+1 10/60)*50
Distance travelled by the train = x*110
When they meet, the two distances are equal
(x+7/6)*50 = 110x
50x+350/6 = 110x
175/3= 60x
x = 175/180 hours
x = (175/180)*60 minutes
x = 175/3 minutes
x = 58 min 20 sec
Therefore they will meet at 9h 58m 20s

Answer by Alan3354(69443)

A taxi left Los Angeles at 7:50 am travelling 50 kph. At 9 am, an express train travelling 110 kph left the same spot travelled by the taxi. At what time did the train overtake the taxi
------------------
Do these the easy way:
In 70 minutes (900 - 750) the taxi went 50*(70/60) km = 175/3 km
The train gains on the taxi at 60 kph (110-50), or 1 km/minute
-----------------
(175/3)/1 = 175/3 minutes = 58 mins 20 seconds
--> 09:58:20 AM